OpenStudy (anonymous):

In simplifying the expression quantity 5 minus 2 end quantity to the second power, Micah followed the steps below. Where did he go wrong? Expression: (5 - 2)2 + 4 • 3 Step 1: 5 - 4 + 4 • 3 Step 2: 5 - 4 + 12 Step 3: 1 + 12 Step 4: 13 Step 1 Step 2 Step 3 Step 4

OpenStudy (anonymous):

Medals Given

OpenStudy (anonymous):

Um, so there seems to me to be a lot of mistakes... Can you write it in the Equation editor for us, please?

OpenStudy (anonymous):

Is this supposed to be your expression? \[ (5-2)^2 +4*3\]

OpenStudy (anonymous):

(5-2) 2+4.3 (3)8/3 your answer is 8

OpenStudy (anonymous):

Never mind i have the answer

OpenStudy (anonymous):

@messiahcollins can you please write the expressions into the Math Editor? I don't understand the question... Given the "word problem", this is how it should look: \[(5-2)^2 = 3^2 = 9\]

OpenStudy (anonymous):

Okay, sorry I couldn't help much!

OpenStudy (anonymous):

what was the first problem she made in pemdas then i found it

OpenStudy (anonymous):

it was step one